OpenStudy (anonymous):

In algebra solve for 6x^4-17x^2+5=0

OpenStudy (anonymous):

6x^4-17x^2+5=0 6x^4-15x^2-2x^2+5=0 3x^2(2x^2-5)-1(2x^2-5)=0 (3x^2-1)(2x^2-5)=0

OpenStudy (anonymous):

so you are saying that x would =(3x^2-1)(2x^2-5)

OpenStudy (anonymous):

no we have to solve dis eqn.

OpenStudy (anonymous):

on solving I think dat we may get x=sqrt(1/3),sqrt(5/2)

OpenStudy (anonymous):

Thank you for you help but I think that the Answer to my question is going to be different than that x would have to equal something that looks like this (3x^2-1)(2x^2-5) not sqrt(1/3), sqrt(5/2)